# Observation of the quantum equivalence principle for matter-waves

### v1

@@ Abstract
Einstein’s general theory of relativity is based on the principle of equivalence — in essence, dating back to Galileo — which asserts that, locally, the effect of a gravitational field is equivalent to that of an accelerating reference frame, so that the local gravitational field is eliminated in a freely falling frame. Einstein’s theory enables this principle to extend to a global description of relativistic space-time, at the expense of allowing space-time to become curved, realising a consistent frame-independent description of nature at the classical level. Einstein’s theory has been confirmed to great accuracy for astrophysical bodies. However, in the quantum domain the equivalence principle involves a gauge phase that is observable if the wavefunction — fundamental to quantum descriptions — allows an object to interfere with itself after being simultaneously at rest in two differently accelerating frames, one being the laboratory (Newtonian) frame and the other in the freely-falling (Einsteinian) frame. This is done with a novel cold-atom interferometer in which one wave packet stays static in the laboratory frame while the other is in free-fall. We follow the relative-phase evolution of the wave packets in the two frames, confirming the equivalence principle in the quantum domain. Our observation is yet another fundamental test of the interface between quantum theory and gravity. The new interferometer also opens the door for further probing of the latter interface, as well as to searches for new physics.
@@ I Introduction
We use in the following the common definition of the equivalence principle, as Einstein originally thought of it: A state of rest in a homogeneous gravitational field is physically indistinguishable from a state of uniform acceleration in a gravity-free space [11]. Let us note that numerous possible extensions of the equivalence principle to the quantum domain (hereafter the “quantum equivalence principle”) have been proposed [12, 13, 14, 15, 16, 17, 18] and here we focus on the most straightforward aspect, which is the quantum phase connecting between the freely-falling frame and the lab frame [19], termed by some the Einsteinian and Newtonian frames, respectively [20]. This phase has a characteristic dependence of $t^{3}$ on the free-fall time $t$[3, 19, 20]. As this phase dependence is based on the assumption that the inertial mass is equal to the gravitational mass, as suggested by Galileo and as confirmed experimentally with early experiments by Newton, Bessel and Eötvös and more recent experiments with improved accuracy [21, 22, 23], we name our unique interferometer the Quantum Galileo Interferometer, or QGI [24].
The prediction that an accelerated object, and specifically a free-falling object acquires a $t^{3}$-phase, has been made quite some time ago by Darwin, Kennard and others [25, 26, 27, 28, 29, 30, 31, 32, 33]. However, it was only recently realized by Marletto and Vedral that it in fact constitutes an extension of the equivalence principle into the quantum domain [19]. While a cubic phase dependence on the interferometer time may also appear in a few standard-practice interferometers [34, 35, 36, 37, 38], observing this free-fall phase in the QGI configuration sampling these two fundamental frames, constitutes a direct test and confirmation of the phase of a free-falling object, as well as a fundamental aspect of the quantum equivalence principle for matter waves.
@@
Consequently, when observing a single wave packet from the two reference frames, the only witness indicating the observer’s frame is a phase that cannot be measured, heralding the indistinguishability of the quantum system between these two frames and thus establishing the equivalence principle in the quantum domain.
Our experiment intended to measure this fundamental phase is motivated by a Gedanken (thought) experiment where at time $t=t_{0}$ the wave packet is split into two paths where in one path it is at rest in the Newtonian frame, due to an applied force that is exactly opposite to gravity, while in the other path it is in free-fall, namely, at rest in the freely-falling Einsteinian frame. To achieve recombination of the two paths, the wave packet in the freely falling path is initially launched upwards by a magnetic-gradient pulse with a velocity $v_{0}$ in the direction opposite to gravity and then falls freely for a time $2T=2v_{0}/g$ (hereafter referred to as the closing condition), after which an identical magnetic-gradient pulse erases the velocity difference between the two wave packets so that the two paths overlap in position and momentum. Fig. 1(A) describes this Gedanken experiment, where the wave packet which is at rest in the lab (Newtonian) frame is termed the “reference wave packet” while the wave packet which is free-falling in the lab (Newtonian) frame is termed the “ballistic wave packet” [39].
@@
For an intuitive understanding of the QGI in the Newtonian frame, we fix the reference wave packet at vertical position $z=0$ so that it does not accumulate any phase, and calculate the phase accumulated by the ballistic wave packet. The latter accumulates no phase in the freely-falling (Einsteinian) frame, and we can use the gauge phase to know its phase in the Newtonian frame. We calculate the phase in the ballistic path in two parts: First, a wave packet starting at $t=0$ in the middle of the interferometer and going forward in time, and then starting from the same point and going backward in time, with these paths ending at $t=\pm T$ and $z=0$ after a free-fall duration of $T$ [Fig. 1(A)]. It follows that the gauge phase at the center of the wave packet for the two parts, and the phase difference between the two end points read
@@
To fully appreciate how the QGI is uniquely suited to measure the quantum equivalence principle, we emphasize the symmetry of the configuration. An observer sitting in the Einsteinian or Newtonian frames will see exactly the same phenomenon, with the roles of the reference wave packet and the ballistic wave packet inverted. This symmetry becomes strikingly clear when one computes the actions in both frames (SM). Assuming that the inertial and gravitational masses are equal and that the acceleration we induce by an external force to form the reference trajectory is exactly opposite to the gravitational acceleration, the action of the ballistic path in the Newtonian frame and that of the reference path in the Einsteinian frame are the same, equal to $-\frac{m}{3}g^{2}T^{3}$[40].
In contrast, in interferometers based on light pulses, there is no such symmetry. There, one must take into account the electromagnetic fields, namely, introduce transformations of Maxwell equations, so that in such interferometers what is in fact verified is the equivalence principle for electromagnetic waves [41, 42] (see SM). In the QGI, no such additional theoretical structure needs to be invoked.
@@ IV Discussion
The cubic phase holds special meaning in the case of our interferometer. In this case it is equal to the gauge phase of the transformation between the Newtonian and Einsteinian frames of reference, and thus constitutes a formulation of an equivalence principle in the quantum domain. Another point of view explaining how the QGI enables to test the quantum equivalence principle lies in the fact that it gives rise to a perfect symmetry between the two frames, as exemplified by the calculation of the actions as seen from both frames. The QGI is perfectly suited for its mission also because it does not require any transformation of electromagnetic waves, so that its physics is only a product of de-Broglie waves.
@@ V Acknowledgments
We thank the BGU nano-fabrication facility for the high-quality chip and the BGU support team, especially Menachem Givon, Zina Binstock and Yaniv Bar-Haim for their support in building and maintaining the experiment. Funding: This work was funded, in part, by the Israel Science Foundation (grants no. 856/18, 1314/19, 3515/20, and 3470/21), and the German-Israeli DIP project (Hybrid devices: FO 703/2-1) supported by the DFG. Additional funding was provided by the Gordon and Betty Moore (GBMF ID 11936), and Simons (MP-TMPS-00005477) foundations. IF thanks an anonymous USA philanthropist, John Moussouris, Jussi Westergren and the Emmy Network for support and research funding. CM and VV thank the Gordon and Betty Moore Foundation and the Templeton Foundation and CM also thanks the Eutopia Foundation, and Ben Vass for supporting her research. W.P.S. is grateful to the Hagler Institute for Advanced Study at Texas A&\&&M University for a Faculty Fellowship, and to Texas A&\&&M AgriLife Research for the support of this work. We thank Yair Margalit, Daniel Rohrlich, Daniel M. Greenberger, Ernst M. Rasel, J. Ströhle and Gary G. Rozenman for helpful discussions.
@@ VII.1 BEC production
Figure 6: Schematic representation of the wires on the atom chip. The thickness of the wires is $2\rm\,\mu m$, and their width ranges from $40-200\rm\,\mu m$. At the center of the chip are five straight wires. The three SG wires have a width of $40\rm\,\mu m$ each, and their centers are separated by $100\rm\,\mu m$ (labeled 5,9,13) with alternating current polarity. The SG wires produce a quadrupole $97\rm\,\mu m$ below the chip surface. The magnetic trap on the chip consists of two straight wires and four U-shaped wires. The straight wires confine the atoms in the z and y axes and have a width of $40\rm\,\mu m$, and $100\rm\,\mu m$ separate their centers (labeled 7,11). The four U-shaped wires confine the atoms along the x-axis and have a width of $200\rm\,\mu m$ each.
@@ VII.9.2 Figure 3
In Fig. 3(A), we show the phase derived from the experimental data against three calculations of $\phi(2T)$: the numerical simulation, the analytical approximation and the expected phase in the Gedanken experiment. We calculate the lines of the analytical approximation and the Gedanken experiment using Eq. 4 and Eq. 19 respectively. The line experimental data was derived by a fit to the data of Fig. 2 (as explained above), and the data points of the numerical simulation were derived as explained in Sec. H. We also plot as a guide to the eye deviations from the $T^{3}$ phase accumulation, by plotting a modified version of the analytical approximation,
@@ VII.10 Magnetic gradient pulses
The area under such a pulse is $I_{\rm max}(\tau+w)$. The scheme has two short kick pulses, and one longer holding pulse. The kick pulses have a rise and fall time of $\tau_{\rm kick}=40\rm\,\mu s$ each, with $w=0$ and $I_{\rm max}=I_{\rm kick}$, so that the area of the pulse is $I_{\rm kick}\cdot\tau_{\rm kick}$. The longer holding pulse has a rise and fall time of $\tau_{\rm hold}=12\rm\,\mu s$, and a duration $w=T_{h}-\tau_{\rm hold}$ of constant current $I_{\rm hold}$ so that the total area of the pulse is $a_{\rm hold}\cdot T_{h}$.
@@
To improve the response of the current source and reduce the effect of the ambient gradient in the chamber, we set the current between the pulses to a finite value, which we call the idle current, and set its value to $I_{\rm idle}=0.47\,\rm mA$. The idle is on during the entire interferometer sequence.. Including the idle current, we get the following requirements for the currents
@@ VII.13 Wave-packet size
Direct experimental measurement of the wave-packet size is a hard task. To begin with, our optical resolution is only $3\,\rm\mu m$. Second, imaging near the chip distorts the image due to diffraction from the chip. Diffraction also originates from the high optical density of the atom cloud. Finally, the magnetic potential of the DKC, as well as the magnetic fields during the SG kicks, are curved as they emanate from the nearby chip wires. These curvatures create lensing (focusing) effects which result in fast changes of the wave-packet size over time. Given these restrictions, when imaging the wave-packet width we observe a width of $3-4\,\rm\mu m$ on the CCD, which indicates that the actual size in most probably below our optical resolution.
@@ VII.14 Connection between gauge phase and equivalence principle
This theoretical summary establishes the connection used in the introduction of the paper between the gauge phase and the equivalence principle. We follow the presentation given in [28, 3] as well as [19]. We will assume, as noted in the introduction, that the inertial mass and gravitational mass are identical. The Lagrangian for a particle with mass $m$ moving in an accelerated frame with the acceleration rate given by $g$ (Einsteinian frame) is:
@@ VII.15 Einstein equivalence principle for Maxwell waves
In KCI the beams splitters and mirrors are formed [41] by an optical two-photon transition in a $\Lambda$-system which is detuned with respect to the upper level. Although the frequency difference of the transition between the ground state and the third level is in the microwave regime, the total momentum transfer determined by the wave vector $k$ of the light fields reads $2\hbar k$, since the light fields are counter propagating. We emphasize that in the KCI, the momentum transfer is always $2\hbar k$ independent of the momentum of the scattered atom.
In contrast, in the QGI the beam splitters and the mirrors are realized by two different mechanisms. Indeed, the linear gravitational potential and the magnetic field gradient during the ballistic motion (holding time) act as mirrors where every momentum is, in principle, perfectly inverted. It is this feature which ensures the symmetry between the reference and the ballistic wave packet, as shown in Table I. The beam splitters in the QGI create the coherent superposition between two internal states of the atom without transferring a substantial momentum. It is the magnetic-gradient pulses which launch and stop the ballistic wave packet.
@@
We conclude by comparing and contrasting this aspect to the QGI. In contrast to the KCI where the atoms have a non-vanishing velocity when they interact with the light fields coherently splitting the atom, in the QGI they have approximately a vanishing velocity during the MW pulses. Indeed, the two magnetic-gradient pulses launching the atom and stopping it, are applied after and before the MW operation on the internal atomic levels. Moreover, the magnetic-gradient pulses, the Fourier-Transform of which may be considered an electromagnetic wave holding a frequency of about 10 kHz, do not require a resonance condition, in contrast to the beam splitters and mirrors in the KCI. Hence, no transformation of these fields is required.

### v2

@@ Abstract
Einstein’s general theory of relativity is based on the principle of equivalence — in essence, dating back to Galileo — which asserts that, locally, the effect of a gravitational field is equivalent to that of an accelerating reference frame, so that the local gravitational field is eliminated in a freely falling frame. Einstein’s theory enables this principle to extend to a global description of relativistic space-time, at the expense of allowing space-time to become curved, realising a consistent frame-independent description of nature at the classical level. Einstein’s theory has been confirmed to great accuracy for astrophysical bodies. However, in the quantum domain the equivalence principle has been predicted to take a unique form involving a gauge phase that is observable if the wavefunction — fundamental to quantum descriptions — allows an object to interfere with itself after being simultaneously at rest in two differently accelerating frames, one being the laboratory (Newtonian) frame and the other in the freely-falling (Einsteinian) frame. To measure this gauge phase we realise a novel cold-atom interferometer in which one wave packet stays static in the laboratory frame while the other is in free-fall. We follow the relative-phase evolution of the wave packets in the two frames, confirming the equivalence principle in the quantum domain. Our observation is yet another fundamental test of the interface between quantum theory and gravity. The new interferometer also opens the door for further probing of the latter interface, as well as to searches for new physics.
@@ I Introduction
We use in the following the common definition of the equivalence principle, as Einstein originally thought of it: A state of rest in a homogeneous gravitational field is physically indistinguishable from a state of uniform acceleration in a gravity-free space [11]. Let us note that numerous possible extensions of the equivalence principle to the quantum domain (hereafter the “quantum equivalence principle”) have been proposed [12, 13, 14, 15, 16, 17, 18, 19] and here we focus on the most straightforward aspect, which is the quantum phase connecting between the freely-falling frame and the lab frame [20], termed by some the Einsteinian and Newtonian frames, respectively [21]. This phase has a characteristic dependence of $t^{3}$ on the free-fall time $t$[3, 20, 21]. As this phase dependence is based on the assumption that the inertial mass is equal to the gravitational mass, as suggested by Galileo and as confirmed experimentally with early experiments by Newton, Bessel and Eötvös and more recent experiments with improved accuracy [22, 23, 24], we name our unique interferometer the Quantum Galileo Interferometer, or QGI [25].
The prediction that an accelerated object, and specifically a free-falling object acquires a $t^{3}$-phase, has been made quite some time ago by Darwin, Kennard and others [26, 27, 28, 29, 30, 31, 32, 33, 34]. While a cubic phase dependence on the interferometer time may also appear in a few standard-practice interferometers [35, 36, 37, 38, 39], observing this free-fall phase in the QGI configuration sampling these two fundamental frames, constitutes a direct test and confirmation of the phase of a free-falling object, as well as a fundamental aspect of the quantum equivalence principle for matter waves [3, 17, 20].
@@
This fundamental phase plays two roles: first, it is the phase difference between the same wave packet as viewed from the Einsteinian and Newtonian frames in a gravitational field (e.g., Eq. 5 in Ref. [21]); second, as explained in Ref. [20], it is the phase between a wave packet in a frame accelerated with acceleration $a=-g$ in an environment devoid of gravity, and an identical wave packet in a frame that is at rest in a gravitational field, such that the two systems are indistinguishable as the phase cannot be measured by any local operation – thus heralding the quantum equivalence principle.
Our experiment intended to measure this fundamental phase — by way of interference with a reference wave packet — follows a Gedanken (thought) experiment where at time $t=t_{0}$ the wave packet is split into two paths where in one path it is at rest in the Newtonian frame, due to an applied force that is exactly opposite to gravity, while in the other path it is in free-fall, namely, at rest in the freely-falling Einsteinian frame. To achieve recombination of the two paths, the wave packet in the freely falling path is initially launched upwards by a magnetic-gradient pulse with a velocity $v_{0}$ in the direction opposite to gravity and then falls freely for a time $2T=2v_{0}/g$ (hereafter referred to as the closing condition), after which an identical magnetic-gradient pulse erases the velocity difference between the two wave packets so that the two paths overlap in position and momentum. Fig. 1(A) describes this Gedanken experiment, where the wave packet that is at rest in the lab (Newtonian) frame is termed the “reference wave packet” while the wave packet that is free-falling in the lab (Newtonian) frame is termed the “ballistic wave packet” [40].
@@
For an intuitive understanding of the QGI in the Newtonian frame, we fix the reference wave packet center at vertical position $z=0$ so that it does not accumulate any propagation phase (beyond the free evolution of the wave packet), and calculate the phase accumulated by the ballistic wave packet. The latter accumulates no propagation phase in the freely-falling (Einsteinian) frame, and we can use the gauge phase to know its phase in the Newtonian frame. We calculate the phase in the ballistic path in two parts: First, a wave packet starting at $t=0$ in the middle of the interferometer and going forward in time, and then starting from the same point and going backward in time, with these paths ending at $t=\pm T$ and $z=0$ after a free-fall duration of $T$ [Fig. 1(A)]. It follows that the gauge phase at the center of the wave packet for the two parts, and the phase difference between the two end points read
@@
In fact, the QGI also demonstrates a quantum equivalence principle between the Newtonian and Einsteinian frames, for a quantum system in a state of spatial superposition. This is due to the complete symmetry of the QGI. An observer sitting in the Einsteinian or Newtonian frames will see exactly the same phenomenon, with the roles of the reference wave packet and the ballistic wave packet inverted. This symmetry becomes strikingly clear when one computes the actions in both frames (SM). Assuming that the inertial and gravitational masses are equal and that the acceleration we induce by an external force to form the reference trajectory is exactly opposite to the gravitational acceleration, the action of the ballistic path in the Newtonian frame and that of the reference path in the Einsteinian frame are the same, equal to $-\frac{m}{3}g^{2}T^{3}$[41].
Finally, the QGI is able to measure the quantum equivalence principle for matter waves, as it utilizes no electromagnetic waves resonant with some atomic transition. In contrast, in interferometers based on light pulses, one must take into account the electromagnetic fields, namely, introduce transformations of Maxwell equations, so that a working interferometer may be described in both frames. Consequently, in such interferometers, what is in fact verified is the equivalence principle for electromagnetic waves [42, 43] (see SM). In the QGI, no such additional theoretical structure needs to be invoked.
@@ IV Discussion
The cubic phase holds special meaning in the case of our interferometer. In this case it is equal to the phase constituting the quantum equivalence principle or the phase of the transformation between the Newtonian and Einsteinian frames of reference. Another point of view explaining how the QGI enables to probe this fundamental phase lies in the fact that it gives rise to a perfect symmetry between the two frames, as exemplified by the calculation of the actions as seen from both frames. The QGI is perfectly suited for its mission also because it does not require any transformation of electromagnetic waves, so that its physics is only a product of de-Broglie waves.
@@ V Acknowledgments
We thank the BGU nano-fabrication facility for the high-quality chip and the BGU support team, especially Menachem Givon, Zina Binstock, Dmitrii Kapusta and Yaniv Bar-Haim for their support in building and maintaining the experiment. Funding: This work was funded, in part, by the Israel Science Foundation (grants no. 856/18, 1314/19, 3515/20, and 3470/21), and the German-Israeli DIP project (Hybrid devices: FO 703/2-1) supported by the DFG. Additional funding was provided by the Gordon and Betty Moore (doi.org/10.37807/GBMF11936), and Simons (MP-TMPS-00005477) foundations. IF thanks an anonymous USA philanthropist, John Moussouris, Jussi Westergren and the Emmy Network for support and research funding. CM and VV thank the Gordon and Betty Moore Foundation and the Templeton Foundation and CM also thanks the Eutopia Foundation, and Ben Vass for supporting her research. W.P.S. is grateful to the Hagler Institute for Advanced Study at Texas A&\&&M University for a Faculty Fellowship, and to Texas A&\&&M AgriLife Research for the support of this work. We thank Časlav Brukner, Daniel Rohrlich, Daniel M. Greenberger, Ernst M. Rasel, Yair Margalit, J. Ströhle, Gary G. Rozenman, Fabio Di Pumpo and Matthias Zimmermann for helpful discussions.
@@ VII.1 BEC production
Figure 6: Schematic representation of the wires on the atom chip. The thickness of the wires is $2\rm\,\mu m$, and their width ranges from $40-200\rm\,\mu m$. At the center of the chip are five straight wires. The three SG wires have a width of $40\rm\,\mu m$ each, and their centers are separated by $100\rm\,\mu m$ (labeled 5, 9, 15) with alternating current polarity. The SG wires produce a quadrupole $97\rm\,\mu m$ below the chip surface. The magnetic trap on the chip consists of two straight wires and four U-shaped wires. The straight wires (labeled 7, 13) confine the atoms in the z and y axes and have a width of $50\rm\,\mu m$, and their centers are separated by $100\rm\,\mu m$. The four U-shaped wires confine the atoms along the x-axis and have a width of $200\rm\,\mu m$ each. This confinement is produced by the inner wires (labeled 1, 19, 21, 39).
@@ VII.9.2 Figure 3
In Fig. 3(A), we show the phase derived from the experimental data against three calculations of $\phi(2T)$: the numerical simulation, the analytical approximation and the expected phase in the Gedanken experiment. We calculate the lines of the analytical approximation and the Gedanken experiment using Eq. 4 and Eq. 19 respectively. The line of experimental data was derived by a fit to the data of Fig. 2 (as explained above), and the data points of the numerical simulation were derived as explained in Sec. H. We also plot as a guide to the eye deviations from the $T^{3}$ phase accumulation, by plotting a modified version of the analytical approximation,
@@ VII.10 Magnetic gradient pulses
The area under such a pulse is $I_{\rm max}(\tau+w)$. The scheme has two short kick pulses, and one longer holding pulse. The kick pulses have a rise and fall time of $\tau_{\rm kick}=40\rm\,\mu s$ each, with $w=0$ and $I_{\rm max}=I_{\rm kick}$, so that the area of the pulse is $I_{\rm kick}\cdot\tau_{\rm kick}$. The longer holding pulse has a rise and fall time of $\tau_{\rm hold}=12\rm\,\mu s$, and a duration $w=T_{h}-\tau_{\rm hold}$ of constant current $I_{\rm hold}$ so that the total area of the pulse is $I_{\rm hold}\cdot T_{h}$.
@@
To improve the response of the current source and reduce the effect of the ambient gradient in the chamber, we set the current between the pulses to a finite value, which we call the idle current, and set its value to $I_{\rm idle}=0.47\,\rm mA$. The idle is on during the entire interferometer sequence. Including the idle current, we get the following requirements for the currents
@@ VII.13 Wave-packet size
Direct experimental measurement of the wave-packet size is a hard task. To begin with, our optical resolution is only $3\,\rm\mu m$. Second, imaging near the chip distorts the image due to diffraction from the chip. Diffraction also originates from the high optical density of the atom cloud. Finally, the magnetic potential of the DKC, as well as the magnetic fields during the SG kicks, are curved as they emanate from the nearby chip wires. These curvatures create lensing (focusing) effects which result in fast changes of the wave-packet size over time. Given these restrictions, when imaging the wave-packet width we observe a width of $3-4\,\rm\mu m$ on the CCD, which indicates that the actual size is most probably below our optical resolution.
@@ VII.14 Connection between gauge phase and equivalence principle
This theoretical summary establishes the connection used in the introduction of the paper between the gauge phase and the equivalence principle. We follow the presentation given in [29, 3] as well as [20]. We will assume, as noted in the introduction, that the inertial mass and gravitational mass are identical. The Lagrangian for a particle with mass $m$ moving in an accelerated frame with the acceleration given by $g$ (Einsteinian frame) is:
@@ VII.15 Einstein equivalence principle for Maxwell waves
In KCI the beam splitters and mirrors are formed [42] by an optical two-photon transition in a $\Lambda$-system which is detuned with respect to the upper level. Although the frequency difference of the transition between the ground state and the third level is in the microwave regime, the total momentum transfer determined by the wave vector $k$ of the light fields reads $2\hbar k$, since the light fields are counter propagating. We emphasize that in the KCI, the momentum transfer is always $2\hbar k$ independent of the momentum of the scattered atom.
In contrast, in the QGI the beam splitters and the mirrors are realized by two different mechanisms. Indeed, the linear gravitational potential and the magnetic field gradient during the ballistic motion (holding time) act as mirrors where every momentum is, in principle, perfectly inverted. It is this feature which ensures the symmetry between the reference and the ballistic wave packet, as shown in Table I. The beam splitters in the QGI ($\pi/2$ MW pulses) create the coherent superposition between two internal states of the atom without transferring a substantial momentum. It is the magnetic-gradient pulses which launch and stop the ballistic wave packet.
@@
We conclude by comparing and contrasting this aspect to the QGI. In the KCI, the interaction with the atoms is very sensitive to a resonance condition, due to a significant Doppler shift ${\bf k}\cdot{\bf v}$. Therefore, changing reference frames, which amounts to changing the velocity of the atoms, must be accompanied by an adequate change of the electromagnetic field frequency, and the transformation of the electromagnetic fields of the beam splitters and mirrors plays an important role in the KCI. In contrast, in the QGI, the magnetic-gradient pulses - the Fourier-Transform of which may be considered as an electromagnetic wave having a frequency of about 10 kHz - do not require a resonance condition to interact efficiently with the atoms. Furthermore, the Rabi frequency of the MW pulses responsible for the internal spin splitting and spin inversion is much larger than the Doppler shift. Hence, no transformation of MW fields between the two frames is required in the QGI interferometer.